# Continual Multi-Robot Learning from Black-Box Visual Place Recognition Models

### v1

@@ Abstract
In the context of visual place recognition (VPR), continual learning (CL) techniques offer significant potential for avoiding catastrophic forgetting when learning new places. However, existing CL methods often focus on knowledge transfer from a known model to a new one, overlooking the existence of unknown black-box models. We explore a novel multi-robot CL approach that enables knowledge transfer from black-box VPR models (teachers), such as those of local robots encountered by traveler robots (students) in unknown environments. Specifically, we introduce Membership Inference Attack, or MIA, the only major privacy attack applicable to black-box models, and leverage it to reconstruct pseudo training sets, which serve as the key knowledge to be exchanged between robots, from black-box VPR models. Furthermore, we aim to overcome the inherently low sampling efficiency of MIA by leveraging insights on place class prediction distribution and un-learned class detection imported from the VPR literature as a prior distribution. We also analyze both the individual effects of these methods and their combined impact. Experimental results demonstrate that our black-box MIA (BB-MIA) approach is remarkably powerful despite its simplicity, significantly enhancing the VPR capability of lower-performing robots through brief communication with other robots. This study contributes to optimizing knowledge sharing between robots in VPR and enhancing autonomy in open-world environments with multi-robot systems that are fault-tolerant and scalable.
@@ I Introduction
Visual place recognition (VPR) enables autonomous robots and self-driving vehicles to classify their place from visual input [1, 2, 3]. While conventional VPR systems rely on supervised learning from direct visual experiences, they face two fundamental limitations: the high cost of collecting training data in new environments and catastrophic forgetting when learning new places. These challenges become particularly acute in long-term autonomous operations where robots must continuously adapt to environmental changes. Recent advancements in continual learning techniques [4]—such as regularization methods, experience replay, and dynamic network expansion—yielded promising results in mitigating catastrophic forgetting and improving continual adaptation performance [5, 6, 7].
This study explores multi-robot continual learning (CL) for visual place recognition (VPR), framing CL as knowledge transfer from an old model to a new one. We extend conventional knowledge transfer between different models of the same robot to knowledge transfer between different models of different robots. This approach offers key advantages, including scalability and fault tolerance. Specifically, a traveling robot can update its knowledge by acquiring information from local robots, avoiding the cost and risk of collecting training data independently. Additionally, sharing VPR knowledge ensures critical information is preserved even if some robots fail. Despite these advantages, multi-robot CL introduces new challenges, such as variations in robot capabilities, communication overhead, and concerns related to privacy and security.
@@
Figure 1: Even simple human-to-human communication or knowledge transfer (KT) allows travelers (students) to avoid getting lost by acquiring VPR knowledge from locals (teachers). In analogy, this study explores a robot-to-robot KT framework. Specifically, we tackle the model inversion problem of reconstructing a (pseudo) training dataset from a black-box teacher with minimal assumptions about the unseen teacher.
One of the key research challenges in extending multi-robot systems is encapsulated in the question: “Can a pseudo-training set be reconstructed from a black-box teacher?” Reconstructing a pseudo-training set from an existing model is a core task in continual learning, as it is essential for integrating new training sets and retraining new models. Many existing approaches rely on the assumption that the existing model is a white box, allowing access to its internal structure or training set. Indeed, in single-robot systems, this assumption often holds, as the existing model belongs to the robot itself. However, in multi-robot systems, especially in open-world multi-robot scenarios where the robots are heterogeneous or unknown robots, making the same assumption is overly optimistic. In fact, reconstructing a training set from a black-box model remains an ongoing area of research in machine learning and is largely an open problem.
In this work, we address the key challenge of reconstructing (pseudo) training samples from black-box teachers using membership inference attacks (MIAs) [8] (Fig. 1). Unlike traditional approaches that analyze the model to reconstruct the training data, MIA generates samples and predicts the likelihood that they were part of the model’s training data. It is the only major privacy attack method that is effective with black-box models. However, this flexibility comes with a downside: low sampling efficiency, particularly when dealing with high-dimensional input data, such as images used in visual place recognition (VPR). The naive approach of randomly generating training sample candidates is highly inefficient. To overcome this limitation, we propose improving the sampling efficiency of MIA by leveraging insights on place class prediction distributions [9] and un-learned class detection [10] from the VPR literature as a prior distribution. We also evaluate the effects of these methods both individually and in combination. Through extensive experiments, we demonstrate that the proposed approach significantly enhances performance in continual learning (CL).
Our key contributions include:
  1. 1.
    This study extends the conventional continual learning (CL) framework from a single-robot to a multi-robot paradigm, facilitating knowledge transfer between robots.
  2. 2.
    It introduces a method for reconstructing pseudo-training samples from a black-box teacher model, tackling challenges in teacher-student VPR knowledge transfer using membership inference attacks (MIA).
  3. 3.
    Experimental results demonstrate that our prior-based approach is remarkably powerful despite its simplicity, significantly enhancing the VPR capability of lower-performing robots through brief communication with other robots.
This study significantly extends our late-breaking paper in [11] and generalizes our previous research on VPR, including continual learning [12], knowledge transfer from teacher to student [9], and knowledge distillation [13], into a novel data-free variant applicable to black-box teachers. As autonomous robots continue to proliferate, universal knowledge transfer protocols and continual domain adaptation will become increasingly important. This research lays the foundation for optimizing robot-to-robot interactions in various VPR scenarios and aims to enhance robotic autonomy in open-world environments.
II Formulation of Multi-Robot CL
@@
This section starts by reviewing the conventional task, which includes the VPR task, supervised learning, and single-robot CL, and builds upon it to formulate multi-robot CL and MIA, which is the focus of our work. Furthermore, as part of the preparation for the evaluation experiments, we introduce a metric for assessing the knowledge transfer cost.
@@
Here, $C$ is a predefined set of place classes, each of which is defined in a real-world region. For instance, in the case of the NCLT public dataset [14] used in our experiments, grid-based partitioning [10] is employed. The workspace is partitioned into $10\times 10$ grid cells in a bird’s-eye coordinate system, with each cell being defined as a place class (Fig. 1). This grid partitioning provides a standard place class; however, as a trade-off, the intra-class variation becomes large, making the classification task more challenging (Fig. 2). One straightforward approach to overcoming the difficulties caused by intra-class variation is optimizing the place definition [15]. This is a fundamental challenge in the VPR community and an ongoing topic of research [3]. While this approach is outside the scope of this paper, we believe it is orthogonal and complementary to the place classification approach presented in this study.
@@
Figure 2: Examples of input images from independent sessions. Each row presents images from different place classes across four sessions, with each column containing four image samples from the corresponding place class. The grid-based partitioning adopted in this paper is the most naive and standard solution to the place class definition issue. However, it results in large intra-class variation, making the VPR task more challenging.
@@
In multi-robot CL, rather than receiving a new training set $T_{0}^{+}$, a new black-box teacher model $M_{0}^{+}$ arrives. In this case, instead of the generic inversion function $I$, a specialized inversion function $I^{*}$ (i.e., MIA) applicable to the black-box teacher model is required to reconstruct the new training data as:
@@
The MIA attempts to reconstruct the training set of a teacher model, which is originally inaccessible, through interactions consisting of queries and responses to the teacher model. The overall process can be expressed as:
@@
Finally, we discuss the knowledge transfer cost. In multi-robot CL, the interaction between the student robot and the black-box teacher occurs in the form of queries and responses. Specifically, the queries from the student robot are automatically generated by the program code, and the code used in this paper consists of short, fixed-size code snippets of just a few dozen lines. On the other hand, the responses from the teacher are pseudo-training samples. These are represented as long real-valued vectors or tensors, which consume communication costs proportional to the number of pseudo-samples. Based on this background, this paper evaluates the knowledge transfer cost based on the number of pseudo-samples sent from the teacher to the student, assuming that the queries from the student to the teacher do not consume communication costs, or that these costs are negligible.
III Dataset Reconstruction using BB-MIA
A key challenge in applying Membership Inference Attacks (MIA) to black-box models, also known as black-box MIA (BB-MIA), is the high-dimensionality of the sample space. In BB-MIA, the attacker (student) cannot directly access the teacher model’s internal parameters or the distribution of its training data. As a result, generating suitable samples to effectively attack the model’s decision boundaries becomes extremely difficult. To address this issue, we approximate the teacher model using a cascade pipeline consisting of two modules: a pre-trained embedding module and a trainable MIA module. The embedding module transforms high-dimensional input images into lower-dimensional embedding vectors, utilizing a scene graph classifier as an intermediate step, which has been validated for its generalizability and effectiveness across various VPR tasks (e.g., [9]). The MIA module incorporates multiple strategies to efficiently sample from the embedding vectors, enabling high-efficiency MIA sampling. Each module is detailed in the following subsections.
III-A Generic Embedding Model
@@
III-B Black-Box Membership Inference Attacks
The problem here is to reconstruct a pseudo-sample set $\{(x,P(y|x))\}$ that approximates the training data of a black-box teacher, given a teacher that takes an embedding vector as input and outputs a class-specific probability map (CPM). This is an inverse problem of supervised learning.
US (Uniform Sampling) Strategy: The simplest strategy involves sampling $x$ of dimension $|C|$ from a uniform distribution, applying L1 normalization, and inputting it into the teacher model to obtain the output CPM $P(y|x)$.
$x_{i}=\frac{{u}_{i}}{\|{u}_{i}\|_{1}},\quad{u}_{i}\sim U,\quad\text{for }i=1,2,\dots,N$ (11)
RR (Reciprocal Rank) Strategy: Although simple, the US strategy is not constrained to approximate the predictive distribution of the classifier model (i.e., scene graph classifier GCN). Therefore, as an alternative to the US strategy, we introduce the reciprocal rank (RR) strategy. As shown in our previous study [9], the output of a classifier model is often approximated by the reciprocal rank feature (RRF), which takes the following form.
$x_{i}=f^{RR}\left(\frac{{u}_{i}}{\|{u}_{i}\|_{1}}\right),\quad{u}_{i}\sim U,\quad\text{for }i=1,2,\dots,N$ (12)
@@ III-B Black-Box Membership Inference Attacks
Entropy Strategy: Although the RR strategy can approximate the predictive distribution of a generice VPR model, it is not constrained to approximate the predictive distribution of the target teacher model. To overcome this issue, we introduce the Entropy strategy. Entropy is frequently used in the self-localization community to find unseen images for VPR [10]. In contrast, we utilize the entropy measure for the novel application of predicting the place classes that were seen during the training phase of the teacher VPR model. The Entropy strategy selects high-quality samples based on the assumption that samples $(x,P(y|x))$ with low entropy of $P(y|x)$ are more likely to be members of the teacher model’s training set. Specifically, we generate an excessive number of samples using the RR strategy and select the top $N$ samples with the lowest entropy score $H$.
$H(x)=-\sum_{i=1}^{|C|}P(y|x_{i},f)\log P(y|x_{i},f).$ (14)
Replay/Prior Strategy: Leaving the category of BB-MIA methods, we also consider the conventional replay-based sampler for benchmarking purposes, which contradicts the assumption of a black-box teacher [20]. This strategy assumes that the teacher is not a complete black-box and retains a subset of the training dataset bundled with the teacher model, and that the student has access to this subset. Such an assumption is relevant in multi-robot systems with superior communication and memory capacities, where the communication channel between the teacher and student is sufficiently wide, and the teacher model’s memory capacity is large enough. As a variant of the standard replay strategy in continual learning, we also introduce a novel strategy called the Prior strategy. Instead of assuming that the teacher model retains its training samples, as in the replay strategy, the Prior strategy assumes that the student model is bundled with its own training samples. These available training samples are then used as additional samples for querying the teacher. Such an assumption is relevant in scenarios where only the student, rather than the teacher, is cooperative in continual learning and where the communication bandwidth is sufficiently large to transmit the training samples.
Mixup Strategy: In addition, we introduce the Mixup strategy, which combines the advantages of the replay strategy and other (RR or Entropy) strategies. This strategy assumes that only a small subset of the training set maintained by the student/teacher model, with a size of $R$ (e.g., $R=1$), is retained. This assumption is relevant when the communication and memory capabilities of a multi-robot system are not as abundant as those envisioned by the Replay strategy, but are still not entirely unavailable. It generates the required $N$ samples by combining the retained $R$ samples with $(N-R)$ samples from the set generated by the RR/Entropy strategy:
@@
This strategy is clearly inapplicable to a black-box teacher and should therefore be used only as an oracle baseline for benchmarking purposes.
IV Experimental Evaluation
We evaluate the performance of the proposed method in a typical multi-robot continual learning (CL) scenario, where a traveler robot (the student) encounters three teachers in succession and receives knowledge transfer via wireless communication. However, there is also the risk of forgetting previously learned place classes. One of our goals is to investigate the trade-off between knowledge acquisition and forgetting.
IV-A Experimental Setup
We used the NCLT dataset [14] in our experiments. The NCLT dataset provides sensor data obtained from a Segway robot navigating across multiple sessions spanning across multiple seasons on the North Campus of the University of Michigan. Specifically, in our VPR tasks, we use images from the robot’s onboard camera as sensor input, annotated with ground-truth viewpoint GPS data. We follow a common protocol for using the NCLT dataset, collected by a single robot, in multi-robot scenarios, where different robots are paired with different sessions [21].

### v2

# Multi-Robot Data-Free Continual Communicative Learning (CCL) from Black-Box Visual Place Recognition Models

@@ Abstract
In emerging multi-robot societies, heterogeneous agents must continually extract and integrate local knowledge from one another through communication, even when their internal models are completely opaque. Existing approaches to continual or collaborative learning for visual place recognition (VPR) largely assume white-box access to model parameters or shared training datasets, which is unrealistic when robots encounter unknown peers in the wild. This paper introduces Continual Communicative Learning (CCL), a data-free multi-robot framework in which a traveler robot (student) continually improves its VPR capability by communicating with black-box teacher models via a constrained query–response channel. We repurpose Membership Inference Attacks (MIA), originally developed as privacy attacks on machine learning models, as a constructive communication primitive to reconstruct pseudo-training sets from black-box VPR teachers without accessing their parameters or raw data. To overcome the intrinsic communication bottleneck caused by the low sampling efficiency of black-box MIA, we propose a prior-based query strategy that leverages the student’s own VPR prior to focus queries on informative regions of the embedding space, thereby reducing the knowledge transfer (KT) cost. Experimental results on a standard multi-session VPR benchmark demonstrate that the proposed CCL framework yields substantial performance gains for low-performing robots under modest communication budgets, highlighting CCL as a promising building block for scalable and fault-tolerant multi-robot systems.
@@ I Introduction
Visual place recognition (VPR) enables autonomous robots and self-driving vehicles to recognize their location from visual input [1, 2, 3, 4, 5]. VPR has been extensively studied from the perspectives of feature representation, robustness to appearance changes, and large-scale deployment [4, 5]. While conventional VPR systems rely on supervised learning from direct visual experiences, they face two fundamental limitations in long-term operation: the high cost of collecting training data in each new environment and catastrophic forgetting when learning new places.
Continual Learning (CL) techniques [6, 7, 8] alleviate catastrophic forgetting by repeatedly adapting a single robot’s model, but they typically assume white-box access to model parameters or replay buffers, which does not hold when robots must collaborate across organizational or vendor boundaries. In future robot-populated societies, diverse agents will coexist in shared environments, each maintaining its own VPR model trained on local experience. A traveler robot should be able to communicate with such local robots and acquire useful knowledge about their environments, even when their internal models and training data remain private or unknown. This motivates a shift from single-robot CL to a broader Continual Communicative Learning (CCL) paradigm, where the primary resource exchanged between agents is not data or parameters but messages over a constrained communication channel.
@@
Figure 1: Conceptual illustration of communicative knowledge transfer (KT). Even simple human-to-human communication allows travelers (students) to avoid getting lost by acquiring place-recognition knowledge from locals (teachers) without seeing their internal mental models. In analogy, this study explores a multi-robot Continual Communicative Learning (CCL) framework, where a robot student interacts with black-box teacher VPR models through a query–response protocol and reconstructs pseudo-training data for continual adaptation.
Communicative learning has recently been proposed as a unified learning formalism that views learning as a bidirectional communication process between teachers and students, subsuming passive learning, active learning, and machine teaching under a single multi-agent framework [9, 10]. It has been instantiated in embodied AI scenarios such as bidirectional human–robot value alignment [11] and communicative navigation with natural gestures [12], highlighting the potential of learning through rich interactions. Our CCL framework can be seen as bringing the same spirit of communicative learning to multi-robot VPR, focusing on continual knowledge transfer among heterogeneous robots.
Figure 1 illustrates this idea. Just as human travelers can avoid getting lost by asking local people for directions without ever seeing their internal mental models, a robot traveler should be able to improve its VPR capability by interacting with local robots through a query–response protocol. The key challenge is that these local robots are often black-box agents: their architectures, training pipelines, and datasets are inaccessible, and only an input–output API is available. A central question for CCL is therefore:
Can a traveler robot reconstruct a pseudo-training set that captures a black-box teacher’s knowledge using communication alone, and use it for continual adaptation?
In this work, we address this question by framing multi-robot continual learning as communicative learning from black-box VPR models. We consider a traveler robot (student) that encounters local robots (teachers) in different environments and aims to extract their place-recognition knowledge without accessing their parameters or data. To this end, we introduce Membership Inference Attacks (MIA) [13]—one of the most widely studied privacy attacks against black-box models [13]—and repurpose them as constructive communication tools. By repeatedly querying a teacher and analyzing its responses, the student reconstructs pseudo-training samples that approximate the teacher’s experience and integrates them into its own model in a CCL fashion.
However, naive use of MIA suffers from extremely low sampling efficiency when dealing with high-dimensional inputs such as images, which translates directly into a prohibitive communication cost. To overcome this limitation, we propose a prior-based query strategy that leverages the student’s own VPR prior—including its prediction distribution and uncertainty estimates—to generate and select informative queries. This strategy focuses communication on regions where the student is uncertain or under-trained, thereby reducing the number of queries required to achieve effective knowledge transfer within CCL.
Our contributions are threefold:
  • We introduce Continual Communicative Learning (CCL) for visual place recognition, formulating multi-robot continual learning from black-box VPR models as a teacher–student knowledge transfer problem.
  • We show that Membership Inference Attacks can be inverted from privacy attacks into data-free communication primitives for reconstructing pseudo-training sets, and we design a prior-based query strategy that significantly improves their sampling efficiency in high-dimensional VPR settings.
  • Through extensive experiments on a standard multi-session VPR benchmark, we demonstrate that the proposed CCL framework significantly improves the performance of low-performing robots under modest communication budgets, suggesting a promising direction for scalable and fault-tolerant multi-robot systems.
II Related Work
II-A Visual Place Recognition
Visual place recognition has been widely studied in robotics and computer vision, with a variety of handcrafted and deep learning-based approaches proposed for robust localization under appearance changes and viewpoint variations [1, 2, 3]. Recent surveys provide comprehensive overviews of deep VPR methods, datasets, and evaluation protocols [4, 5]. Unlike conventional VPR methods that assume direct access to raw images and labels, our work focuses on CCL from black-box VPR models, where only an input–output API is available and communication cost is explicitly modeled.
II-B Continual Learning and Knowledge Transfer
Continual Learning (CL) addresses catastrophic forgetting when a model is trained on a sequence of tasks or data distributions [6]. Recent surveys summarize CL theory, methods, and applications across vision, language, and reinforcement learning [7, 8]. Most CL methods, including replay-based and distillation-based approaches [14, 15, 16], assume white-box access to model parameters or stored exemplars. In contrast, our CCL framework targets a multi-robot scenario where a student must perform continual adaptation by communicating with black-box teachers through pseudo-samples reconstructed via MIA.
II-C Membership Inference and Model Inversion
Membership Inference Attacks (MIA) aim to infer whether a given sample was part of a model’s training set, raising fundamental privacy concerns for machine learning systems [13]. Recent surveys provide taxonomies of MIA methods and defenses for both white-box and black-box models [13]. While most prior work views MIA as a threat, some emerging studies explore data-free knowledge transfer and model inversion using pseudo-samples generated from black-box models [17, 18]. Our work pushes this line further by explicitly treating black-box MIA as a communicative primitive within CCL for multi-robot VPR.
II-D Communicative Learning and Multi-Agent Communication
Communicative learning has been proposed as a unified formalism that models learning as communication between teachers and students, with internal “minds” that reason about each other’s beliefs and intentions [9, 10]. This paradigm has been instantiated in embodied settings such as in situ bidirectional human–robot value alignment [11] and communicative navigation with natural gestures [12]. Our CCL framework can be viewed as a concrete realization of communicative learning in multi-robot VPR, focusing on continual knowledge transfer among heterogeneous VPR models.
In multi-agent reinforcement learning (MARL), communication learning has been extensively studied as a means to improve coordination and scalability [19, 20]. Surveys on multi-agent deep reinforcement learning with communication summarize architectures and protocols for differentiable communication channels [20]. Our work is complementary: rather than optimizing communication for joint control, we treat communication itself as a vehicle for continual knowledge transfer in CCL, with an explicit focus on black-box VPR models and data-free pseudo-samples.
III Formulation of Multi-Robot CCL for VPR
@@
This section starts by reviewing the conventional task, which includes the VPR task, supervised learning, and single-robot CL, and builds upon it to formulate multi-robot CCL and black-box MIA, which are the focus of our work. Furthermore, as part of the preparation for the evaluation experiments, we introduce a metric for assessing the communicative knowledge transfer cost.
@@
Here, $C$ is a predefined set of place classes, each of which is defined in a real-world region. For instance, in the case of the NCLT public dataset [21] used in our experiments, grid-based partitioning [22] is employed. The workspace is partitioned into $10\times 10$ grid cells in a bird’s-eye coordinate system, with each cell being defined as a place class (Fig. 1). This grid partitioning provides a standard place class; however, as a trade-off, the intra-class variation becomes large, making the classification task more challenging (Fig. 2). One straightforward approach to overcoming the difficulties caused by intra-class variation is optimizing the place definition [23]. This is a fundamental challenge in the VPR community and an ongoing topic of research [3]. While this approach is outside the scope of this paper, it is orthogonal and complementary to the CCL-based place classification approach presented in this study.
@@
Figure 2: Examples of input images from independent sessions. Each row presents images from different place classes across four sessions, with each column containing four image samples from the corresponding place class. The grid-based partitioning adopted in this paper is a standard solution to the place class definition issue. However, it results in large intra-class variation, making the CCL-based VPR task more challenging.
@@
In multi-robot CCL, rather than receiving a new training set $T_{0}^{+}$, a new black-box teacher model $M_{0}^{+}$ arrives. In this case, instead of the generic inversion function $I$, a specialized inversion function $I^{*}$ (i.e., black-box MIA) applicable to the teacher model is required to reconstruct the new training data as:
@@
The MIA attempts to reconstruct the training set of a teacher model, which is originally inaccessible, through communicative interactions consisting of queries and responses to the teacher model. The overall CCL process can be expressed as:
@@
Finally, we discuss the communicative knowledge transfer cost. In multi-robot CCL, the interaction between the student robot and the black-box teacher occurs in the form of queries and responses. Specifically, the queries from the student robot are automatically generated by the program code, and the code used in this paper consists of short, fixed-size code snippets of just a few dozen lines. On the other hand, the responses from the teacher are pseudo-training samples. These are represented as long real-valued vectors or tensors, which consume communication costs proportional to the number of pseudo-samples. Based on this background, this paper evaluates the knowledge transfer cost based on the number of pseudo-samples sent from the teacher to the student, assuming that the queries from the student to the teacher do not consume communication costs, or that these costs are negligible.
IV Dataset Reconstruction for CCL using Black-Box MIA
A key challenge in applying Membership Inference Attacks (MIA) to black-box models, also known as black-box MIA (BB-MIA), is the high-dimensionality of the sample space [13]. In BB-MIA, the attacker (student) cannot directly access the teacher model’s internal parameters or the distribution of its training data. As a result, generating suitable samples to effectively probe the model’s decision boundaries becomes extremely difficult.
To address this issue in the context of CCL for VPR, we approximate the teacher model using a cascade pipeline consisting of two modules: a pre-trained embedding module and a trainable MIA module. The embedding module transforms high-dimensional input images into lower-dimensional embedding vectors, utilizing a scene graph classifier as an intermediate step, which has been validated for its generalizability and effectiveness across various VPR tasks (e.g., [24]). The MIA module incorporates multiple strategies to efficiently sample from the embedding vectors, enabling high-efficiency BB-MIA sampling. Each module is detailed in the following subsections.
IV-A Generic Embedding Model
@@
IV-B Black-Box Membership Inference Attacks
The problem here is to reconstruct a pseudo-sample set $\{(x,P(y\mid x))\}$ that approximates the training data of a black-box teacher, given a teacher that takes an embedding vector as input and outputs a class-specific probability map (CPM). This is an inverse problem of supervised learning and serves as the core communicative mechanism in our CCL framework.
US (Uniform Sampling) Strategy: The simplest strategy involves sampling $x$ of dimension $|C|$ from a uniform distribution, applying L1 normalization, and inputting it into the teacher model to obtain the output CPM $P(y\mid x)$:
$x_{i}=\frac{{u}_{i}}{\|{u}_{i}\|_{1}},\quad{u}_{i}\sim U,\quad\text{for }i=1,2,\dots,N.$ (11)
RR (Reciprocal Rank) Strategy: Although simple, the US strategy is not constrained to approximate the predictive distribution of the classifier model (i.e., scene graph classifier GCN). Therefore, as an alternative to the US strategy, we introduce the reciprocal rank (RR) strategy. As shown in our previous study [24], the output of a classifier model is often approximated by the reciprocal rank feature (RRF), which takes the following form:
$x_{i}=f^{RR}\left(\frac{{u}_{i}}{\|{u}_{i}\|_{1}}\right),\quad{u}_{i}\sim U,\quad\text{for }i=1,2,\dots,N,$ (12)
@@ IV-B Black-Box Membership Inference Attacks
Entropy Strategy: Although the RR strategy can approximate the predictive distribution of a generic VPR model, it is not constrained to approximate the predictive distribution of the target teacher model. To overcome this issue, we introduce the Entropy strategy. Entropy is frequently used in the self-localization community to find unseen images for VPR [22]. In contrast, we utilize the entropy measure for the novel application of predicting the place classes that were seen during the training phase of the teacher VPR model. The Entropy strategy selects high-quality samples based on the assumption that samples $(x,P(y\mid x))$ with low entropy of $P(y\mid x)$ are more likely to be members of the teacher model’s training set. Specifically, we generate an excessive number of samples using the RR strategy and select the top $N$ samples with the lowest entropy score $H$:
$H(x)=-\sum_{i=1}^{|C|}P(y\mid x_{i},f)\log P(y\mid x_{i},f).$ (14)
Replay/Prior Strategy: Leaving the category of BB-MIA methods, we also consider the conventional replay-based sampler for benchmarking purposes, which contradicts the assumption of a purely black-box teacher [14]. This strategy assumes that the teacher is not a complete black-box and retains a subset of the training dataset bundled with the teacher model, and that the student has access to this subset. Such an assumption is relevant in multi-robot systems with superior communication and memory capacities, where the communication channel between the teacher and student is sufficiently wide, and the teacher model’s memory capacity is large enough.
As a variant of the standard replay strategy in continual learning, we also introduce a novel strategy called the Prior strategy. Instead of assuming that the teacher model retains its training samples, as in the Replay strategy, the Prior strategy assumes that the student model is bundled with its own training samples. These available training samples are then used as additional samples for querying the teacher. Such an assumption is relevant in scenarios where only the student, rather than the teacher, is cooperative in CCL and where the communication bandwidth is sufficiently large to transmit the training samples.
Mixup Strategy: In addition, we introduce the Mixup strategy, which combines the advantages of the replay strategy and other (RR or Entropy) strategies. This strategy assumes that only a small subset of the training set maintained by the student/teacher model, with a size of $R$ (e.g., $R=1$), is retained. This assumption is relevant when the communication and memory capabilities of a multi-robot system are not as abundant as those envisioned by the Replay strategy, but are still not entirely unavailable. It generates the required $N$ samples by combining the retained $R$ samples with (NR)(N-R) samples from the set generated by the RR/Entropy strategy:
@@
This strategy is clearly inapplicable to a strict black-box teacher and should therefore be used only as an oracle baseline for benchmarking purposes in our CCL setting.
V Experimental Evaluation of CCL for Multi-Robot VPR
We evaluate the performance of the proposed CCL framework in a typical multi-robot continual learning scenario, where a traveler robot (the student) encounters three teachers in succession and receives communicative knowledge transfer via wireless communication. However, there is also the risk of forgetting previously learned place classes. One of our goals is to investigate the trade-off between knowledge acquisition and forgetting under realistic communication budgets.
V-A Experimental Setup
We used the NCLT dataset [21] in our experiments. The NCLT dataset provides sensor data obtained from a Segway robot navigating across multiple sessions spanning multiple seasons on the North Campus of the University of Michigan. Specifically, in our VPR tasks, we use images from the robot’s onboard camera as sensor input, annotated with ground-truth viewpoint GPS data. We follow a common protocol for using the NCLT dataset, collected by a single robot, in multi-robot scenarios, where different robots are paired with different sessions [28].